# KOALA++: Efficient Kalman-Based Optimization of Neural Networks with Gradient-Covariance Products

### v1

@@ 3.3 Least-Squares Estimate of $P_{k-2}$
In the first variant of our approach, we solve the least-squares problem without imposing any structural constraints on the covariance estimate. Given the relation $v_{k-1}=H_{k-1}P_{k-2}$, we compute $P_{k-2}\in\mathbb{R}^{n\times n}$ by introducing Lagrange multipliers $\lambda\in\mathbb{R}^{1\times n}$ and write the Lagrangian
$\mathcal{L}(P_{k-2},\lambda)=\|P_{k-2}\|_{F}^{2}+(H_{k-1}P_{k-2}-v_{k-1})\lambda^{T}.$ (11)
This formulation admits a closed-form solution [^1] See Appendix B.1 for a detailed derivation., which is:
$P_{k-2}=\frac{H_{k-1}^{T}v_{k-1}}{||H_{k-1}||^{2}}.$ (12)
@@
$H_{k}P_{k-2}=\alpha_{k}v_{k-1},\quad\text{where}\quad\alpha_{k}:=\frac{H_{k}H_{k-1}^{T}}{\|H_{k-1}\|^{2}}.$ (13)
Finally, we can get the vanilla $v_{k}$ update as follows
@@
To improve stability and theoretical consistency with classical Kalman filtering, we introduce a second variant that enforces symmetry in the estimation of $P_{k-2}$. In particular, the additional regularization term $P_{k-2}=P_{k-2}^{T}$ is added to the objective. As in the previous derivation, we obtain a closed-form solution for $P_{k-2}$ [^2] See Appendix B.2 for a detailed derivation.
@@
$w_{k}$ $=\frac{(H_{k}v_{k-1}^{T})(H_{k-1}H_{k-1}^{T})-(H_{k-1}v_{k-1}^{T})(H_{k}H_{k-1}^{T})}{\|H_{k-1}\|^{4}}.$ (16)
To simplify notation and implementation, we adopt a unified update rule for both variants. In the unconstrained case, we simply set $w_{k}=0$, effectively reducing the update to the asymmetric version. This allows the general form of the update to be written compactly as
$\boxed{v_{k}=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q+w_{k}H_{k-1}.}$ (17)
To maintain consistency with the original KOALA formulation and to simplify computation in the early stages of training, we initialize the directional covariance product $v_{1}=H_{1}P_{0}$ using a scaled identity assumption on the initial covariance matrix $P_{0}$. Specifically, we set:
@@
     For simplicity, denote $H_{k}=\nabla L_{k}(\theta_{k-1})$Calculate $\alpha_{k},\lambda_{k},w_{k}$ respectively from Equations (13) (8) (16)
@@
$v_{k}$ $=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q+w_{k}H_{k-1}$ (19)
$\theta_{k}$ $=\theta_{k-1}-\frac{\eta_{k}L_{k}(\theta_{k-1})}{H_{k}v_{k}^{T}+H_{k}QH_{k}^{T}+R}\cdot(v_{k}^{T}+QH_{k}^{T})$ (20)
@@ 4 Experiments
We evaluate KOALA++ on a range of vision and language tasks to demonstrate its generality, stability, and convergence behavior compared to existing optimizers. Our experiments are organized into three parts: image classification, language modeling and ablation studies. Due to space constraints, we report a representative subset of results in the main text. Additional evaluations, including ablations, hyperparameter tuning, and extended visualizations, are provided in Appendix D.
@@ 4.1 Image Classification
We follow the experimental setup of the original KOALA paper davtyan2022koala for CIFAR-10 and CIFAR-100 classification tasks, including data augmentation, model architectures, and optimization settings. We introduce two modifications: (i) training is repeated with three random seeds (42, 3407, and 2025) to ensure statistical robustness, and (ii) we adopt a two-stage learning rate decay strategy at the 100th and 150th epochs for 200-epoch training (inspired by successful multi-step decay strategies reported in prior works such as rice2020overfitting ), which improves generalization compared to the single-step decay used in the original paper. The results are shown in Table  2.
Table 2: Experimental results on CIFAR-10 and CIFAR-100 with different optimizers for 100 and 200 epochs. Best and second best performance is indicated in boldface and with a blue shading.
@@
Adam 7.09 0.30 6.09 0.29
KOALA-M 6.34 0.30 5.85 0.25
[…]
Adam 6.43 0.27 5.74 0.33
KOALA-M 8.00 0.24 8.16 0.33
@@
W-ResNet-50-2 SGD 23.83 6.35 22.47 5.96
Adam 23.73 6.64 24.07 7.06
@@
(1) CIFAR-100 + ResNet-50 + Advanced Optimizers. To evaluate KOALA++ in a standardized yet competitive setting, we follow the benchmarking protocol introduced in KOALA davtyan2022koala , and compare our method with a wide range of recent optimizers on the CIFAR-100 dataset using ResNet-50, trained for 100 epochs. All optimizers are tested under the same learning rate schedule (StepLR) and training configuration. For each optimizer, we adopt hyperparameters as suggested in the original papers or official implementations. The settings and results are shown in Table 3
@@
Yogi reddi2018adaptive 33.99 10.90 lr=$10^{-2}$, $\beta_{1}$=0.9, $\beta_{2}$=0.999, $\epsilon$=$10^{-3}$
@@
Fromage bernstein2020distance 24.65 6.71 lr=$10^{-2}$ (recommended in GitHub [^3] https://github.com/jxbz/fromage#voulez-vous-du-fromage)
@@
[…]
@@
Figure 1 presents the training and validation dynamics on CIFAR-100 with ResNet-50 under a multi-step learning rate schedule at epochs 30, 60, and 90. Compared to SGD, Adam, AdaFisher, and KOALA-M, our proposed KOALA++ demonstrates consistently lower validation loss in log-scale, along with improved generalization as shown by the lower Top-1 error.
@@
Moreover, KOALA++ maintains competitive or superior stability across all training phases, without exhibiting the overfitting spikes or stagnation observed in other optimizers. These observations collectively validate KOALA++ as a robust and efficient optimizer for large-scale vision tasks.
@@
As shown in Table 4, KOALA++ performs on par with AdaFisher across a range of architectures, while exhibiting consistently lower standard deviations, indicating greater stability across random seeds. All KOALA++ results were obtained without extensive hyperparameter adjustment, highlighting the robustness of the method in the default settings. It is also worth noting that on Swin-Tiny, KOALA++ reaches strong performance with significantly smaller variance, demonstrating better adaptability to Transformer-based architectures.
@@ 4.2 Language Model
We evaluate KOALA++ on the Wikitext-2 dataset, following the same language modeling benchmark setting introduced in the AdaFisher paper gomes2024adafisher . The model is a scaled-down version of GPT-1 with four masked self-attention layers and approximately 28 million parameters. We adopt the same architecture, tokenizer, and 50-epoch training schedule as used in the original benchmark.
In line with the original AdaFisher benchmark (See Github [^6] https://github.com/AtlasAnalyticsLab/AdaFisher/tree/main), we report the test perplexity (PPL) corresponding to the best-performing model on the validation set, along with the total training time, to provide a holistic view of performance and efficiency. The results are summarized in Table 5. In addition to reporting test perplexity and total training time, we further provide visualizations to better illustrate the training dynamics(See Appendix D)
@@ 4.3 Ablation Study
Due to space constraints, we report only the core ablation comparing KOALA-M and two variants of KOALA++ on CIFAR-100 with ResNet-50. The default KOALA++ uses a symmetric gain approximation, while KOALA++ (NS) adopts an asymmetric covariance. As shown in Figures 2, KOALA++ achieves faster convergence and better final performance. KOALA++ (NS) is competitive but less stable in later stages, underscoring the benefit of structural symmetry. Full ablations are provided in Appendix C.
@@ 5 Conclusions, Limitations and Future Work
We proposed KOALA++, a principled and efficient optimization framework that extends the Kalman-based KOALA algorithm by incorporating a lightweight directional covariance approximation. KOALA++ retains the stability and simplicity of its predecessor while substantially improving its capacity to capture structured uncertainty in high-dimensional loss landscapes. Through extensive experiments across vision and language tasks, we demonstrate that KOALA++ achieves competitive or superior performance compared to both first- and second-order optimizers, while maintaining training efficiency on par with standard first-order methods.
Limitations. A notable limitation of our method is that the approximated covariance matrix $P_{k}$ is not explicitly constrained to be positive semi-definite (PSD). Although we empirically observe that the eigenvalues of $P_{k}$ remain stable and bounded throughout training (see Appendix B.3), the lack of a formal PSD guarantee may pose risks in certain settings or applications involving sensitive numerical stability. Addressing this limitation efficiently is left to future work.
Future Work. To address this issue, future work may explore enforcing PSD constraints via projection, spectral regularization, or reparameterization techniques. Additionally, we plan to investigate the integration of KOALA++ with adaptive gradient clipping and apply it to large-scale pretraining scenarios such as vision-language models, diffusion models or instruction-tuned LLMs.

### v2

@@ 3.3 Least-Squares Estimate of $P_{k-2}$
In the first variant of our approach, we solve the least-squares problem without imposing any structural constraints on the covariance estimate. Given the relation $v_{k-1}=H_{k-1}P_{k-2}$, we compute $P_{k-2}\in\mathbb{R}^{n\times n}$ by introducing the Lagrange multipliers $\mu\in\mathbb{R}^{1\times n}$ and write the Lagrangian
$\mathcal{L}(P_{k-2},\lambda)=\|P_{k-2}\|_{F}^{2}+(H_{k-1}P_{k-2}-v_{k-1})\mu^{\top}.$ (11)
This formulation admits the closed-form solution [^1] See Appendix B.1 for a detailed derivation.
$P_{k-2}=\frac{H_{k-1}^{\top}v_{k-1}}{\|H_{k-1}\|^{2}}.$ (12)
@@
$H_{k}P_{k-2}=\alpha_{k}v_{k-1},\quad\text{where}\quad\alpha_{k}:=\frac{H_{k}H_{k-1}^{\top}}{\|H_{k-1}\|^{2}}.$ (13)
Finally, we can obtain the following vanilla $v_{k}$ update
@@
To improve stability and theoretical consistency with classical Kalman filtering, we introduce a second variant that enforces symmetry in the approximation of $P_{k-2}$. In particular, the additional regularization term $P_{k-2}=P_{k-2}^{\top}$ is added to the objective. As in the previous derivation, we obtain a closed-form solution for $P_{k-2}$ [^2] See Appendix B.2 for a detailed derivation.
@@
$r_{k}$ $=\frac{(H_{k}v_{k-1}^{\top})(H_{k-1}H_{k-1}^{\top})-(H_{k-1}v_{k-1}^{\top})(H_{k}H_{k-1}^{\top})}{\|H_{k-1}\|^{4}}.$ (16)
To simplify notation and implementation, we adopt a unified update rule for both variants. In the unconstrained case, we simply set $r_{k}=0$, effectively reducing the update to the asymmetric version. This allows the general form of the update to be written compactly as
$\boxed{v_{k}=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q+r_{k}H_{k-1}.}$ (17)
To maintain consistency with the original KOALA formulation and to simplify the calculation in the early stages of training, we initialize the directional covariance product $v_{1}=H_{1}P_{0}$ using a scaled identity assumption on the initial covariance matrix $P_{0}$. Specifically, we set:
@@
  For simplicity, denote $H_{k}=\nabla L_{k}^{\top}(\theta_{k-1})$Calculate $\alpha_{k},\lambda_{k},r_{k}$ respectively from Equations (13), (8), and (16)
@@
$v_{k}$ $=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q+r_{k}H_{k-1}$ (19)
$\theta_{k}$ $=\theta_{k-1}-\frac{\eta_{k}L_{k}(\theta_{k-1})}{H_{k}v_{k}^{\top}+H_{k}QH_{k}^{\top}+R}\cdot(v_{k}^{\top}+QH_{k}^{\top})$ (20)
@@ 4 Experiments
We evaluate KOALA++ on a range of vision and language tasks to demonstrate its generality, stability, and convergence behavior compared to existing optimizers. Our experiments are organized into three parts: image classification, language modeling and ablation studies. Due to space constraints, we report a representative subset of results in the main text. Additional evaluations, including ablations, hyperparameter tuning, and extended visualizations, are provided in Appendix E.
@@ 4.1 Image Classification
We follow the experimental setup of the original KOALA paper davtyan2022koala for CIFAR-10 and CIFAR-100 classification tasks, including data augmentation, model architectures, and optimization settings. We introduce two modifications for CIFAR-10: (i) training is repeated with three random seeds (42, 3407, and 2025) to ensure statistical robustness, and (ii) we adopt a two-stage learning rate decay strategy at the 100th and 150th epochs for 200-epoch training (inspired by successful multi-step decay strategies reported in prior works such as rice2020overfitting ), which improves generalization compared to the single-step decay used in the original paper. The results [^3] Some of the earlier reported results were obtained from existing public benchmarks, and therefore mean and standard deviation values are not available for all entries. To ensure reproducibility, we reran the experiments under identical settings; the full code and configurations are available at: https://github.com/Sumxiaa/KOALA_Plus_Plus. are shown in Table  2.
Table 2: Experimental results on CIFAR-10 and CIFAR-100 with different optimizers for 100 and 200 epochs. Best (blue) and second best (orange) performances are in boldface and with shading.
@@
Adam 6.960.06 0.270.05 6.350.13 0.270.04
KOALA-M 6.290.04 0.370.02 6.060.14 0.280.06
@@
[…]
KOALA-M 7.620.13 0.300.04 7.670.43 0.320.06
[…]
Adam 24.250.34 7.060.15 24.140.12 7.470.34
[…]
W-ResNet-50-2 SGD 23.030.29 5.970.11 23.500.15 6.470.16
Adam 23.990.17 6.690.07 23.740.19 7.100.17
@@
(1) CIFAR-100 + ResNet-50 + Advanced Optimizers. To evaluate KOALA++ in a standardized yet competitive setting, we follow the benchmarking protocol introduced in KOALA davtyan2022koala , and compare our method with a wide range of recent optimizers on the CIFAR-100 dataset using ResNet-50, trained for 100 epochs. All optimizers are tested under the same learning rate schedule (StepLR) and training configuration. For each optimizer, we adopt hyperparameters as suggested in the original papers or official implementations. The settings and results are shown in Table 3.
@@
[…]
@@
[…]
@@
KOALA-M 23.71 6.58 Redid the experiments with our implementation settings
@@
Figure 1 presents the training and validation dynamics on CIFAR-100 with ResNet-50 under a multi-step learning rate schedule at epochs 30, 60, and 90. Compared to SGD, Adam, AdaFisher, and KOALA-M, our proposed KOALA++ demonstrates consistently lower validation loss on the logarithmic scale, along with improved generalization as shown by the lower Top-1 error.
@@
Moreover, KOALA++ maintains competitive or superior stability throughout all training phases, without exhibiting the overfitting spikes or stagnation observed in other optimizers. These observations collectively validate KOALA++ as a robust and efficient optimizer for large-scale vision tasks.
@@
As shown in Table 4, KOALA++ performs on par with AdaFisher across a range of architectures, while exhibiting consistently lower standard deviations, indicating greater stability among random seeds. All KOALA++ results were obtained without extensive hyperparameter adjustment, highlighting the robustness of the method in the default settings. It is also worth noting that on Swin-Tiny, KOALA++ reaches strong performance with significantly smaller variance, demonstrating better adaptability to Transformer-based architectures.
Efficiency profiling.
Beyond accuracy, we also compare the computational efficiency of KOALA++ with strong baselines. KOALA++ achieves runtime and memory comparable to Adam(W), while being substantially more efficient than Shampoo and AdaHessian. A detailed comparison, including per-epoch time, FLOPs, and peak memory usage on ResNet-50 and Swin-Tiny, is provided in the Appendix E.3.
@@ 4.2 Language Model
We evaluated KOALA++ on the Wikitext-2 dataset, following the same language modeling benchmark setting introduced in the AdaFisher paper gomes2024adafisher . The model is a scaled-down version of GPT-1 with four masked self-attention layers and approximately 28 million parameters. We adopt the same architecture, tokenizer, and 50-epoch training schedule as used in the original benchmark.
In line with the original AdaFisher benchmark (see Github [^7] https://github.com/AtlasAnalyticsLab/AdaFisher/tree/main), we report the test perplexity (PPL) corresponding to the best-performing model on the validation set, along with the total training time, to provide a holistic view of performance and efficiency. The results are summarized in Table 5. In addition to reporting test perplexity and total training time, we further provide visualizations to better illustrate the training dynamics (see Appendix E).
@@ 4.3 Ablation Study
Due to space constraints, we report only the core ablation comparing KOALA-M and two variants of KOALA++ on CIFAR-100 with ResNet-50. The default KOALA++ uses a symmetric gain approximation, while KOALA++ (NS) adopts an asymmetric covariance. As shown in Figure 2, KOALA++ achieves faster convergence and better final performance. KOALA++ (NS) is competitive but less stable in later stages, underscoring the benefit of structural symmetry. Full ablations are provided in the Appendix D.
@@ 5 Conclusions, Limitations and Future Work
We proposed KOALA++, an efficient optimization method that extends the Kalman-based KOALA algorithm by incorporating a lightweight directional covariance approximation. KOALA++ retains the stability and simplicity of its predecessor, while substantially improving its capacity to capture general structured uncertainty in high-dimensional loss landscapes. Through extensive experiments across vision and language tasks, we demonstrate that KOALA++ achieves competitive or superior performance compared to both first- and second-order optimizers, while maintaining training efficiency on par with standard first-order methods.
Limitations. A notable limitation of our method is that the approximated covariance matrix $P_{k}$ is not explicitly constrained to be positive semi-definite (PSD). Although we empirically observe that the eigenvalues of $P_{k}$ remain stable and bounded throughout the training (see the Appendix B.3), the lack of a formal PSD guarantee may pose risks in certain settings or applications involving sensitive numerical stability. Addressing this limitation efficiently is left to future work.
Future Work. To address this issue, future work may explore enforcing PSD constraints via projection, spectral regularization, or reparameterization techniques. Additionally, we plan to investigate the integration of KOALA++ with adaptive gradient clipping and apply it to large-scale pretraining scenarios such as vision-language models, diffusion models, or instruction-tuned LLMs.
@@ Acknowledgements
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state our main contributions and support them with theoretical and experimental results (see Sections 1 , 3, and 4).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations in Section  5, including assumptions on covariance symmetry and potential instability.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The assumptions are made in Section  3 and the derivations are provided in the Appendix B.1 B.2
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All experimental settings are described in Section 4 and Appendix E, including models, datasets, and hyperparameters.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We plan to release the code upon publication, and all datasets used (CIFAR-10/100, WikiText-2) are publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Details of model architectures, optimizer configurations, learning rates, and schedules are provided in Section  4 and Appendix E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report averages over 3–5 seeds and include standard deviations in Table 4 and Appendix D.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe compute resources in Appendix E, including training time on an H100 GPU.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We adhere to the NeurIPS Code of Ethics. No private or sensitive data is used.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: The paper does not discuss societal impacts, as it focuses solely on the design and evaluation of optimization algorithms, which are technical contributions without direct societal applications.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not describe such safeguards, as it does not involve the release of high-risk models or datasets such as pretrained language models, image generators, or scraped data.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All external assets used in the paper (e.g., the AdaFisher benchmark) are properly credited, and their licenses and terms of use (e.g., MIT License) have been respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The reimplementation of the KOALA++ optimizer is a new asset introduced in this paper. It is well documented, and usage instructions will be provided alongside the code repository.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The study does not involve human participants, so IRB approval and risk disclosures are not applicable.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We used LLMs for proofreading and structuring (e.g., ChatGPT), but they were not used as part of the core methodology in this research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.